OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zepdrix

ganeshie8 (ganeshie8):

do u knw the 3 forms for a line :- 1) point-slope form 2) slope-intercept form 3) standard form ?

OpenStudy (anonymous):

Yes.

ganeshie8 (ganeshie8):

then its easy :) for part 1, simply write the point-slope form of line : point = (3, -2) slope = 2

ganeshie8 (ganeshie8):

whats the point-slope form of a line ?

OpenStudy (anonymous):

2x-y-8?

ganeshie8 (ganeshie8):

nope, point slope of form is : y - y1 = m(x - x1)

ganeshie8 (ganeshie8):

plugin (x1, y1) = (3, -2) m = 2

ganeshie8 (ganeshie8):

you wud get :- y - (-2) = 2(x - 3)

ganeshie8 (ganeshie8):

thats it, we're done wid part1

OpenStudy (anonymous):

Okay.

ganeshie8 (ganeshie8):

for part2, do u knw slope-intercept form ?

OpenStudy (anonymous):

yes.

ganeshie8 (ganeshie8):

wat is it

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

So.

OpenStudy (anonymous):

It would be

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

rearrange the equation form part1

ganeshie8 (ganeshie8):

y - (-2) = 2(x - 3) y + 2 = 2x - 6 y = 2x - 8

ganeshie8 (ganeshie8):

we're done wid part2 also

ganeshie8 (ganeshie8):

wat about part3 ?

OpenStudy (anonymous):

Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

yes, whats the standard form ?

ganeshie8 (ganeshie8):

can u check ur notes / google a bit and tell me whats the "standard form of line" ?

OpenStudy (anonymous):

It would be

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

y = 2 (x-4)?

ganeshie8 (ganeshie8):

nope. standard form of line is : Ax + By = C

ganeshie8 (ganeshie8):

from part2, we have equation in slope-intercept form as : y = 2x - 8

ganeshie8 (ganeshie8):

u need to change it to standard form

ganeshie8 (ganeshie8):

y = 2x - 8 8 = 2x - y 2x - y = 8

ganeshie8 (ganeshie8):

thats the standard form okay ?

OpenStudy (anonymous):

-2x + y = -8

OpenStudy (anonymous):

Would this be it?

ganeshie8 (ganeshie8):

in standard form, coefficent of x must always be POSITIVE. so, multiply ur equation by -1

ganeshie8 (ganeshie8):

-2x + y = -8 is same as 2x - y = 8

OpenStudy (anonymous):

Okay, so I got it?

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

just show ur final equation as :- 2x - y = 8

OpenStudy (anonymous):

Okay, thank you!

ganeshie8 (ganeshie8):

np :)